20 gm of NaOH is present in 250 ml of H2O .Calculate  the active mass of the solution.

Grade:12

2 Answers

Vikas TU
7 years ago
Active mass is the molarity itself in other terms as well.
Therfore the moles of NaOH => 20/40 => 0.5  mol
Molarity = 0.5*1000/250 => 2M
This would be the active mass of the solution.
arpan chettri
5 years ago
We know,
Active mass=(wt/mw)×(1000/v)
Wt=weight of solute
Mw=molecular weight
V=volume of solution in ml
 
Active mass= (20/40)×(1000/250)
                      =0.5×4
                      =2M